OpenStudy (anonymous):

In the polar area of the zoo, there are 3 polar bears, 12 penguins, and 5 reindeer. write the ratio of each part to the total @iGreen.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

polar bears = 3/20 penguins = 12/20 = 3/5 reindeer = 5/20 = 1/4 is thst right

OpenStudy (anonymous):

Yes! That's perfect, good job!

OpenStudy (anonymous):

@Nnesha